OpenStudy (anonymous):

Factorize x^4-12x^2-64=0

mathslover (mathslover):

64 = 8^2 right?

OpenStudy (anonymous):

middle term breaking

OpenStudy (anonymous):

breaking 12x^2

OpenStudy (anonymous):

(x^2-16)(x^2+4)

mathslover (mathslover):

ok .. : \[\large{x^4+4x^2-16x^2-64=0}\] right?

OpenStudy (anonymous):

absolutely right

OpenStudy (anonymous):

then factorize it again

mathslover (mathslover):

ok now .. \[\large{x^2(x^2+4)-16x^2-64=0}\] right?

OpenStudy (anonymous):

(x^2-16)(x^2+4)

OpenStudy (anonymous):

(x^2-16)(x^2+4)= (x-4)(x+4)(x^2+4)

mathslover (mathslover):

right : \[\large{(a^2-b^2)=(a+b)(a-b)}\]

mathslover (mathslover):

\[\large{x^2-4^2=(x+4)(x-4)}\]

OpenStudy (anonymous):

x=+4,-4

OpenStudy (anonymous):

the formula is right but i need to solve this without formula

OpenStudy (anonymous):

hoe do you know that????? x=+4,-4

OpenStudy (anonymous):

x2=16 squaring x=+4,-4

OpenStudy (anonymous):

then after that the hard part

OpenStudy (anonymous):

using x+4=0

OpenStudy (anonymous):

sorry x2+4=0

OpenStudy (anonymous):

x^2=-4 squaring x=+2i,-2i

OpenStudy (anonymous):

you mean x^2+4=0

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

its done ok fine thanks for help every body